The following bond investment transactions were completed during a recent year by Starks Company: Year 1...

  1. The following bond investment transactions were completed during a recent year by Starks Company:

    Year 1
    Jan. 31 Purchased 36, $1,000 government bonds at 100 plus accrued interest of $180 (one month). The bonds pay 6% annual interest on July 1 and January 1.
    July 1 Received semiannual interest on bond investment.
    Aug. 30 Sold 15, $1,000 bonds at 98 plus $150 accrued interest (two months).

    a. Journalize the entries for these transactions.

    Assume a 360-day year. Do not round interim calculations. Round final answers to nearest dollar. For a compound transaction, if an amount box does not require an entry, leave it blank.

    Year 1, Jan. 31
    • Cash
    • Interest Expense
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
    • Cash
    • Interest Receivable
    • Interest Revenue
    • Loss on Sale of Investments
    • Notes Receivable
    • Cash
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
    Year 1, July 1
    • Cash
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
    • Cash
    • Interest Payable
    • Interest Receivable
    • Investments-Government Bonds
    • Notes Receivable
    • Cash
    • Gain on Sale of Investments
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
    Year 1, Aug. 30
    • Cash
    • Gain on Sale of Investments
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Gain on Sale of Investments
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Loss on Sale of Investments
    • Cash
    • Gain on Sale of Investments
    • Interest Payable
    • Interest Receivable
    • Interest Revenue
    • Cash
    • Gain on Sale of Investments
    • Interest Receivable
    • Investments-Government Bonds
    • Loss on Sale of Investments

    Feedback

    b. Provide the December 31, Year 1, adjusting journal entry for semiannual interest earned on the bonds.

    Assume a 360-day year. Do not round interim calculations. Round final answers to nearest dollar.

    Year 1, Dec. 31
    • Cash
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
    • Cash
    • Interest Receivable
    • Interest Revenue
    • Investments-Government Bonds
    • Notes Receivable
Answer #1
a
Year 1, Jan. 31 Investments-Government Bonds 36000
Interest Receivable 180
       Cash 36180
Year 1, July 1 Cash 1080 =36000*6%*6/12
       Interest Receivable 180
      Interest Revenue 900 =36000*6%*5/12
Year 1, Aug. 30 Cash 14850 =(15000*0.98)+150
Loss on Sale of Investments 300
       Interest Revenue 150 =15000*6%*2/12
      Investments-Government Bonds 15000 =15*1000
b
Year 1, Dec. 31 Interest Receivable 630 =(36-15)*1000*6%*6/12
       Interest Revenue 630